OpenStudy (anonymous):

Find all zeros of f(x)=x^3-8 x^2+18 x-12. I know to take the first and last coefficients (in this case 1 and 12) and take their factors (1 having only a factor of 1 and 12 having 1, 2, 3, 4, 6, 12). I then use synthetic division to find a zero. I have found 2 to be a zero. Then I take the new quadratic which is: x^2-8x+18 and input it into the quadratic formula. So I got these as zeros: 2, 4+i√2, 4-i√2 My math site for school kicked me back as wrong. Where is my error?

OpenStudy (xapproachesinfinity):

eh okay so you did the Descartes rule of sign perhaps you made an error

OpenStudy (xapproachesinfinity):

redo the process and be more careful

OpenStudy (xapproachesinfinity):

According to wolfram there are 3 real zero no complex one

OpenStudy (anonymous):

Yeah I've done it twice, I'm about to try it again.

OpenStudy (anonymous):

Does wolfram explain the algebra of the problem?

OpenStudy (xapproachesinfinity):

No you have to pay to get the details, but you don't need that you need to work it out! it is just about errors in your steps

OpenStudy (xapproachesinfinity):

http://www.wolframalpha.com/input/?i=f(x)%3dx%5e3-8+x%5e2%2b18+x-12 this is the graph

OpenStudy (xapproachesinfinity):

i think you have got one zero correct 2 is right but find the other two

OpenStudy (anonymous):

Thank you. I'm up to the synthetic division and so far I'm good. Still only getting 2.

OpenStudy (xapproachesinfinity):

that's good! stop there! find the two remaining roots

OpenStudy (anonymous):

well I found one error. used the dang initial polynomial numbers instead of the x^-6x+6 that I would get after the SD. But that can't be factored so I have to use the quadratic formula.

OpenStudy (xapproachesinfinity):

Yeah! you need quadratic to find the other zeros

OpenStudy (xapproachesinfinity):

welcome! did you mean 3 plus or minus root(3)

OpenStudy (anonymous):

Yes I did.

OpenStudy (xapproachesinfinity):

okay good!

OpenStudy (xapproachesinfinity):

I think there is an alternative way when it comes to degree 3 poly

OpenStudy (xapproachesinfinity):

\(\large \rm f(x)=x^3-8x^2+18x-12=x^3-2x^2-6x^2+18x-12\\=x^2(x-2)-6(x^2-3x+2)=x^2(x-2)-6(x-2)(x-1)\)

OpenStudy (anonymous):

So what is happening there?

OpenStudy (anonymous):

factor by grouping?

OpenStudy (xapproachesinfinity):

you factor x-2 and you will get the other part that you get by SD yes! this by grouping (but adding one step) i play with one term to make it factorable

OpenStudy (xapproachesinfinity):

it takes some practice to use this technique, but it doesn't work every time lol but i guarantee that it will work for degree 3 or even degree 4 but 5 up things get crazy

OpenStudy (anonymous):

I see now. oh I like it this way much better, but it won't always work. Which is why we are using the quadratic formula.

OpenStudy (xapproachesinfinity):

one other way i go with this is test some number like f(1) f(-1) f(2) simple cases to know if they are potential zeros

OpenStudy (anonymous):

So plug and solve using the factors of the first and last coefficients right?

OpenStudy (xapproachesinfinity):

no it will work all the time for degree 3

OpenStudy (xapproachesinfinity):

i plug in a potential zero and try to use that to factor

OpenStudy (xapproachesinfinity):

but i would say that SD or LD are what you need to think of more than this

OpenStudy (anonymous):

ehh long division, I love it. /sarcasm

OpenStudy (xapproachesinfinity):

yeah whenever something looks hard to factor go for SD or LD but make sure you are doing it carefully! or otherwise you will be thrown off